# Adaptive Softassign via Hadamard-Equipped Sinkhorn

### v3

@@ 3.2 Softassign Transition
$\mathcal{P}_{sk}({X})={X}\circ{SK}^{({X})}={X}\circ(\mathbf{r}^{T}\otimes\mathbf{c})$ (21)
@@
$\mathcal{P}_{sk}(X)=\mathcal{P}_{sk}(X\circ(\mathbf{u}^{T}\otimes\mathbf{v})).$ (22)
@@ Appendix C Proof of Proposition 3
$\mathcal{P}_{sk}({X})={X}\circ{SK}^{({X})}={X}\circ(\mathbf{r}^{T}\otimes\mathbf{c})$ (42)
@@
$\mathcal{P}_{sk}({X})$ $={D}_{(\mathbf{r})}{X}{D}_{(\mathbf{c})}$ (43)
$={X}\circ\underbrace{(\mathbf{r}^{T}\otimes\mathbf{c})}_{{SK}^{({X})}}.$
@@ Appendix D Proofs of Lemma 1, Lemma 2, and Lemma 3
$\mathcal{P}_{sk}(X)=\mathcal{P}_{sk}(X\circ(\mathbf{u}^{T}\otimes\mathbf{v})).$ (44)
@@
$\mathcal{P}_{sk}(Y)=Y\circ(\mathbf{r}_{Y}^{T}\otimes\mathbf{c}_{Y}).$ (45)
@@
$\mathcal{P}_{sk}(X\circ(\mathbf{u}^{T}\otimes\mathbf{v}))$ $=X\circ(\mathbf{u}^{T}\otimes\mathbf{v})\circ(\mathbf{r}_{Y}^{T}\otimes\mathbf{c}_{Y})$ (46)
$=X\circ((\underbrace{\mathbf{u}\circ\mathbf{r}_{Y}}_{\mathbf{r}_{1}})^{T}\otimes(\underbrace{\mathbf{v}\circ\mathbf{c}_{Y}}_{\mathbf{c}_{1}}))$
$=X\circ(\mathbf{r}_{1}^{T}\otimes\mathbf{c}_{1})$
@@
Since $X\circ(\mathbf{r}_{1}^{T}\otimes\mathbf{c}_{1})$ is a doubly stochastic matrix, $\mathbf{r}_{1}^{T}\otimes\mathbf{c}_{1}={SK^{(X)}}$ according to the Proposition 3.
@@ Appendix F Baselines and visualization of experiments
Updates of Reference
The reference [32], Dynamic softassign and adaptive parameter tuning for graph matching, has been enhanced and renamed to CSGO: Constrained-Softassign Gradient Optimization for Large Graph Matching.

### v4

@@ 3.2 Softassign Transition
$\mathcal{P}_{sk}({X})={X}\circ{SK}^{({X})}={X}\circ(\mathbf{r}\otimes\mathbf{c}^{T})$ (21)
@@
$\mathcal{P}_{sk}(X)=\mathcal{P}_{sk}(X\circ(\mathbf{u}\otimes\mathbf{v}^{T})).$ (22)
@@ Appendix C Proof of Proposition 3
$\mathcal{P}_{sk}({X})={X}\circ{SK}^{({X})}={X}\circ(\mathbf{r}\otimes\mathbf{c}^{T})$ (42)
@@
In elementwise form, for each entry $(i,j)$ of the resulting matrix, we have:
$\left[D_{(\mathbf{r})}XD_{(\mathbf{c})}\right]_{ij}=r_{i}\cdot X_{ij}\cdot c_{j}$
We now observe that $\mathbf{r}\otimes\mathbf{c}^{T}$ is a matrix with entries $(\mathbf{r}\otimes\mathbf{c}^{T})_{ij}=r_{i}c_{j}$. Hence, we can express the product as:
$(D_{(\mathbf{r})}XD_{(\mathbf{c})})_{ij}=X_{ij}\cdot(\mathbf{r}\otimes\mathbf{c}^{T})_{ij}$
This is precisely the Hadamard (elementwise) product of $X$ with the matrix $\mathbf{r}\otimes\mathbf{c}^{T}$:
$D_{(\mathbf{r})}XD_{(\mathbf{c})}=X\circ(\mathbf{r}\otimes\mathbf{c}^{T})$
Thus, we conclude:
$\mathcal{P}_{\mathrm{sk}}(X)=D_{(\mathbf{r})}XD_{(\mathbf{c})}=X\circ(\mathbf{r}\otimes\mathbf{c}^{T}).$
@@ Appendix D Proofs of Lemma 1, Lemma 2, and Lemma 3
$\mathcal{P}_{sk}(X)=\mathcal{P}_{sk}(X\circ(\mathbf{u}\otimes\mathbf{v}^{T})).$ (43)
@@
$\mathcal{P}_{sk}(Y)=Y\circ(\mathbf{r}_{Y}\otimes\mathbf{c}_{Y}^{T}).$ (44)
@@
$\mathcal{P}_{sk}(X\circ(\mathbf{u}\otimes\mathbf{v}))$ $=X\circ(\mathbf{u}\otimes\mathbf{v}^{T})\circ(\mathbf{r}_{Y}\otimes\mathbf{c}_{Y}^{T})$ (45)
$=X\circ((\underbrace{\mathbf{u}\circ\mathbf{r}_{Y}}_{\mathbf{r}_{1}})\otimes(\underbrace{\mathbf{v}\circ\mathbf{c}_{Y}}_{\mathbf{c}_{1}})^{T})$
$=X\circ(\mathbf{r}_{1}\otimes\mathbf{c}_{1}^{T})$
@@
Since $X\circ(\mathbf{r}_{1}\otimes\mathbf{c}_{1}^{T})$ is a doubly stochastic matrix, $\mathbf{r}_{1}\otimes\mathbf{c}_{1}^{T}={SK^{(X)}}$ according to the Proposition 3.
@@ Appendix F Baselines and visualization of experiments
Updates
Version 2. The reference [32], Dynamic softassign and adaptive parameter tuning for graph matching, has been enhanced and renamed to CSGO: Constrained-Softassign Gradient Optimization for Large Graph Matching.
Version 3. Correct the typo in the expression of the Sinkhorn projection by replacing
$\mathcal{P}_{\mathrm{sk}}(X)=X\circ(\mathbf{r}^{T}\otimes\mathbf{c})$ (58)
with the correct formulation
$\mathcal{P}_{\mathrm{sk}}(X)=X\circ(\mathbf{r}\otimes\mathbf{c}^{T}).$ (59)